Solve: 1/2 sec(x)=1/3 on [0,2π].
Answer by stanbon(75887)

Solve: 1/2 sec(x)=1/3 on [0,2π].
------
sec(x) = (1/3)/(1/2) = 2/3
------
There is no answer as sec(x) <= -1 or sec(x) >= 1
=====================================================
Cheers,
Stan H.
=====================